Q1) Given the utility function:
U = X3/4 .Y1/4
Find out the optimal quantities of the two commodities using Lagrange method and simplified procedure, if it is given that price of X is Rs.6 and price of Y is Rs.3 and income is equal to Rs.120.
1
Expert's answer
2020-08-17T10:08:56-0400
"MU_x=\frac {3}{4}(\frac {y}{x})^{\frac {1}{4}}"

"MU_y=\frac {1}{4}(\frac {x}{y})^{\frac {3}{4}}"

"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"

"y=1.5x"

"2x+1.5x=40"


"x=11.4"

"y=17.1"